A ball is projected vertically upwards with velocity v if another ball in the same manner thrown (i.e. with v and vertically upwards ) after 'n' second n can be ant positive number
1) find time after wich ball will meet (if possible)
2) hieght from projection plane at which they will meet?
1
Expert's answer
2018-09-25T10:22:09-0400
1) For the first ball:
h=v(t+n)-(g(t+n)^2)/2
For the second ball:
h=vt-(gt^2)/2
So,
v(t+n)-(g(t+n)^2)/2=vt-(gt^2)/2
vn=g/2(n^2+2tn)
v=g/2(n+2t)
t=(2v/g-n)/2=v/g-n/2.
2)
h=v(v/g-n/2)-g/2 (v/g-n/2)^2
h=v^2/2g-(gn^2)/8